OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −(x + 6) y − 4 = 2(x + 6) y + 6 = −(x − 4) y + 6 = 2(x − 4)

OpenStudy (angela210793):

(x-x1)/(x2-x1)=(y-y1)/(y2-y1) (x+6)/8=(y-4)/-4

OpenStudy (anonymous):

u got the wrong options ....... thats what i get.. y - 4 = 1 (x +6) --- 2

OpenStudy (angela210793):

i guess u can do the rest

OpenStudy (anonymous):

check again ... angela210793

OpenStudy (angela210793):

i checked it....and it loos ok..... (If it isn't can u pls tell me whr did i go wrong?) :)

OpenStudy (anonymous):

see m = (y' -y) --------- (x'-x) and th eq is (y - y")=m(x-x")

OpenStudy (angela210793):

yea i know tht kind of eq...but my eq is also fine....

OpenStudy (angela210793):

u found slope 1st i didn't....

OpenStudy (anonymous):

yea but chk tht -4...

OpenStudy (angela210793):

y2-y1=0-4=-4 isn't it? O.o

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

sry i did it wrong

OpenStudy (angela210793):

it's ok :):)

OpenStudy (anonymous):

but the option aren't right?